### v1

# On a mean-field Pontryagin minimum principle for stochastic optimal control

Abstract.
This papers outlines a novel extension of the classical Pontryagin minimum (maximum) principle to stochastic optimal control problems. Contrary to the well-known stochastic Pontryagin minimum principle involving forward-backward stochastic differential equations, the proposed formulation is deterministic and of mean-field type. The Hamiltonian structure of the proposed Pontryagin minimum principle is achieved via the introduction of an appropriate gauge variable. The gauge freedom can be used to decouple the forward and reverse time equations; hence simplifying the solution of the underlying boundary value problem. We also consider infinite horizon discounted cost optimal control problems. In this case, the mean-field formulation allows converting the computation of the desired optimal control law into solving a pair of forward mean-field ordinary differential equations. The proposed mean-field formulation of the Pontryagin minimum principle is tested numerically for a controlled inverted pendulum and a controlled Lorenz-63 system.
1. Introduction
We consider the optimal control problem for a controlled stochastic differential equation (SDE) of the form
(1) ${\rm d}X_{t}=b(X_{t}){\rm d}t+G(X_{t})U_{t}{\rm d}t+\Sigma^{1/2}{\rm d}B_{t},\qquad X_{0}=a,$
under either the finite horizon cost function
(2) $J_{T}(a,U_{0:T})=\mathbb{E}\left[\int_{0}^{T}\left(c(X_{t})+\frac{1}{2}U_{t}^{\rm T}R^{-1}U_{t}\right){\rm d}t+f(X_{T})\right]$
or the infinite horizon discounted cost function
(3) $J(a,U_{0:\infty})=\mathbb{E}\left[\int_{0}^{\infty}e^{-\gamma t}\left(c(X_{t})+\frac{1}{2}U_{t}^{\rm T}R^{-1}U_{t}\right){\rm d}t\right].$
Here $B_{t}$ denotes $d_{x}$-dimensional Brownian motion, $\Sigma\in\mathbb{R}^{d_{x}\times d_{x}}$ the symmetric positive definite diffusion matrix, $R\in\mathbb{R}^{d_{u}\times d_{u}}$ a symmetric positive definite weight matrix, $G(x)\in\mathbb{R}^{d_{x}\times d_{u}}$ the possibly position-dependent control matrix, $c(x)\geq 0$ the running cost, $f(x)\geq 0$ the terminal cost, and $\gamma>0$ the discount factor. See, for example, [2] for more details. We also introduce the weighted norm $\|\cdot\|_{R}$ via
(4) $\|u\|^{2}_{R}=u^{\rm T}R^{-1}u.$
The aim is to find the closed loop control law $U_{t}=u_{t}(X_{t})$ that minimizes the cost $J_{T}$ over the set of admissible control laws.
It is well known, that the associated optimal control problems for ordinary differential equations (ODEs); i.e. $\Sigma=0$ in (1), can be solved by the Pontryagin minimum (maximum) principle (PMP) [21]. The PMP has been extended to SDEs (1) in terms of forward-backward SDEs [20, 2]. In this paper, we provide an alternative extension via a deterministic mean-field formulation, which leads to interacting particle implementations in the form of Hamiltonian ODEs. The formulation is attractive from a computational perspective since it allows for a decoupling of the forward and backward evolution equations for finite horizon optimal control problems and a forward simulation for infinite horizon optimal control problems. The proposed mean-field approach also delivers closed loop control laws in contrast to the classical PMP for ODEs [21]. The proposed methodology can then also be used to approximate viscosity solutions for controlled ODEs by setting $\Sigma=\epsilon I$ for sufficiently small $\epsilon$ [3].
Mean-field formulations and their interacting particle approximations have been considered before in the context of the Hamilton–Jacobi–Bellman (HJB) equation [2]. The key idea is to connect the time evolution of the value function to a probability density function evolving under an appropriate nonlinear Fokker–Planck equation and its mean-field representation [14, 10, 25, 11]. In this paper, we directly approximate the gradient of the value function and derive appropriate Hamiltonian evolution equations; thus extending mean-field formulations to the PMP.
The remainder of the paper is structured as follows. The following Section 2 summarizes key mathematical results for finite horizon optimal control problems at the level of the underlying HJB and Fokker–Planck equations [2, 19]. The novel mean-field formulation of the PMP is provided in Section 3 starting from an Eulerian variational perspective in terms of the HJB equation. Following [27, 9], the arising action functional is subsequently transformed into a Lagrangian formulation with trajectories following mean-field Hamiltonian ODEs. The infinite horizon optimal control problem is treated in Section 4. Here the key result is a pair of forward mean-field ODEs, which asymptotically deliver the desired closed loop control law. A number of numerical approximation techniques are discussed in Section 5; including an approximation based on discrete Schrödinger bridges [16, 28] as outlined in Section 5.2. The mean-field formulation is tested for two infinite horizon discounted cost (3) optimal control problems in Section 6. The first example is provided by a controlled inverted pendulum [18] while the second example is provided by a controlled Lorenz-63 system [12].
2. Mathematical background
It is well-known [3, 2] that, assuming sufficient regularity, the desired closed loop control law is provided by
(5) $u_{t}(x)=-RG(x)^{\rm T}\nabla_{x}v_{t}(x)$
with the optimal value function $v_{t}(x)$ satisfying the HJB equation
(6) $-\partial_{t}v_{t}=b\cdot\nabla_{x}v_{t}+\frac{1}{2}\Sigma:D_{x}^{2}v_{t}+c-\frac{1}{2}\|RG(x)^{\rm T}\nabla_{x}v_{t}\|^{2}_{R},\qquad v_{T}=f.$
Here $A:B=\sum_{i,j}a_{ij}b_{ij}$ [19]. In case of deterministic dynamics; i.e. $\Sigma=0$ in (1), the PMP provides the open loop optimal control $U_{t}=-RG(X_{t})^{\rm T}P_{t}$ along the solution $X_{t}$ of (1) by solving the ODEs
(7a) $\dot{X}_{t}$ $=b(X_{t})-G(X_{t})RG(X_{t})^{\rm T}P_{t},$
(7b) $-\dot{P}_{t}$ $=(D_{x}b(X_{t}))^{\rm T}P_{t}-\frac{1}{2}\nabla_{x}\|RG^{\rm T}(X_{t})P_{t}\|_{R}^{2}+\nabla_{x}c(X_{t})$
in $(X_{t},P_{t})$ subject to the boundary conditions $X_{0}=a$ and $P_{T}=\nabla_{x}f(X_{T})$ [21]. Formulation (7) offers significant computational savings compared to (6) and has been widely used especially in the context of nonlinear model predictive control (NMPC) [7, 23]. The PMP for ODEs has been extended to SDEs (1) giving rise to the, so called, stochastic PMP with (7) replaced by a pair of forward-backward SDEs [2].
In this paper, we develop a Hamiltonian mean-field formulation in the spirit of the PMP-based Hamiltonian ODE formulation (7), which, in turn, can be approximated by interacting particle systems. Since the PMP directly targets the gradient of the value function $v_{t}(x)$; i.e., $P_{t}=y_{t}(X_{t})$ with $y_{t}(x)=\nabla_{x}v_{t}(x)$, we note that
(8) $-\partial_{t}y_{t}=\nabla_{x}(y_{t}\cdot b)+\frac{1}{2}\nabla_{x}(\Sigma:D_{x}y_{t})+\nabla_{x}c-\frac{1}{2}\nabla_{x}\|RG^{\rm T}y_{t}\|^{2}_{R},\quad y_{T}=\nabla_{x}f,$
which follows form (6) assuming sufficient regularity of the value function.
Substituting the optimal control law (5) into (1), the law $\pi_{t}$ of the states $X_{t}$ satisfies the Fokker–Plank equation
(9) $\partial\pi_{t}=-\nabla_{x}\cdot\left(\pi_{t}(b+Gu_{t})\right)+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\pi_{t}),\qquad\pi_{0}=p_{0},$
for given initial density $p_{0}$ [19].
We assume throughout this paper that the diffusion matrix $\Sigma$ is position independent. However, the proposed PMP extends to position dependent $\Sigma(x)$. We also allow for $\Sigma=0$, in which case our formulation provides the optimal closed loop control law instead of the open loop solution provided by (7). In this case, adding diffusion via $\Sigma=\epsilon I$ with $\epsilon\to 0$ sufficiently small allows one to numerically approximate viscosity solutions of (6) [3].
3. Pontryagin mean-field formulation
In this section, we develop a novel mean-field approach to the PMP starting from an Eulerian perspective in the spirit of ideal fluid dynamics [27]. Specifically, we combine the HJB equation (6) and a Liouville equation for the flow of densities $\rho_{t}(x)$ under a single variational principle. In order to do so, we assume that the initial states $X_{0}$ are distributed according to a given distribution $p_{0}(x)$. Following [27], we then transform the variational principle into a Lagrangian framework in terms of evolving positions $X_{t}\in\mathbb{R}^{d_{x}}$ and momenta $P_{t}\in\mathbb{R}^{d_{x}}$. These equations are Hamiltonian and provide the novel PMP formulation. While we consider the finite horizon optimal control problem (2) in this section; an extension to infinite horizon optimal control problems can be found in Section 4.
3.1. Eulerian formulation
In this subsection, we develop a variational formulation for the HJB equation (6). The first step is to formally decouple the diffusion term in (6) through an additional function $\phi_{t}$, i.e.,
(10a) $-\partial_{t}v_{t}$ $=b\cdot\nabla_{x}v_{t}+\frac{1}{2}\Sigma:D_{x}^{2}\phi_{t}+c-\frac{1}{2}\|RG(x)^{\rm T}\nabla_{x}v_{t}\|^{2}_{R},$
(10b) 00 0 $=\phi_{t}-v_{t}.$
The second step is to introduce Lagrange multipliers $\rho_{t}(x)$ and $\beta_{t}(x)$ in order to enforce (10) in terms of the action functional
(11a) $\mathcal{S}(v,\rho,\beta)$ $=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\partial_{t}v_{t}+\nabla_{x}v_{t}\cdot b+c-\frac{1}{2}\|RG^{\rm T}\nabla_{x}v_{t}\|^{2}_{R}\right\}\rho_{t}{\rm d}x\,{\rm d}t\,+$
(11b) $\qquad\qquad\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\beta_{t}\cdot\nabla_{x}(\phi_{t}-v_{t})+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\phi_{t})\,\right\}\rho_{t}{\rm d}x\,{\rm d}t\,+$
(11c) $\qquad\qquad\int_{\mathbb{R}^{d_{x}}}\left\{(v_{T}-f)\,\rho_{T}-v_{0}p_{0}\right\}{\rm d}x,$
where we have used $\Sigma:D_{x}^{2}v=\nabla_{x}\cdot(\Sigma\nabla_{x}v)$.
More specifically, the Lagrange multiplier $\beta_{t}$ enforces the constraint (10b) (up to a constant), while the Lagrange multiplier $\rho_{t}$ enforces the HJB formulation (10a). Furthermore, variation with respect to $v_{t}$ leads, after integration by parts, to the Liouville equation
(12) $\partial_{t}\rho_{t}=-\nabla_{x}\cdot\left(\rho_{t}(b-\beta_{t}-GRG^{\rm T}\nabla_{x}v_{t})\right),\qquad\rho_{0}=p_{0},$
in the density $\rho_{t}$. The associated ODE is given by
(13) $\dot{X}_{t}=b(X_{t})-\beta_{t}(X_{t})-G(X_{t})RG(X_{t})^{\rm T}\nabla_{x}v_{t}(X_{t}),\qquad X_{0}\sim p_{0}.$
While (10a) is obtained independently of the choice of $\beta_{t}$, a further variation with respect to $\phi_{t}$ leads to
(14) $-\nabla_{x}\cdot(\rho_{t}\beta_{t})+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\rho_{t})=0$
and, hence, to
(15) $\beta_{t}=\frac{1}{2}\Sigma\nabla_{x}\log\rho_{t}.$
With this choice of $\beta_{t}$, the associated Liouville equation (12) turns into the Fokker–Planck equation (9) for the SDE (1). In other words, (13) and (15) lead to an ODE mean-field reformulation of the SDE (1) as considered previously, for example, in [15].
3.2. Lagrangian formulation
Let us now convert (11) into a corresponding Lagrangian formulation [27]. Let $X_{\tau}(a)$ denotes the solution of (13) at time $\tau=t$ with initial condition $X_{0}=a$. Here we have introduced $\tau$ as a new time variable in order to distinguish the two time derivatives arising from the Lagrangian (total time-derivative) and Eulerian (partial time-derivative) perspective. Furthermore, the initial conditions $a\sim p_{0}$ may be viewed as labels in the sense of Lagrangian fluid dynamics. We therefore consider functions $x(a)$, $p(x)$, $u(a)$, and $\beta(a)$ and note that
(16) $\int_{\mathbb{R}^{d_{x}}}g(X_{\tau}(a))\,p_{0}(a){\rm d}a=\int_{\mathbb{R}^{d_{x}}}g(x)\,\rho_{t}(x){\rm d}x$
for any function $g(x)$. Next we convert (11) into its corresponding Lagrangian formulation and consider in particular
(17a) $\mathcal{I}$ $:=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\partial_{t}v_{t}(x)\,\rho_{t}(x){\rm d}x\,{\rm d}t$
(17b) $=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\partial_{t}v_{t}(X_{\tau=t}(a))\,p_{0}(a){\rm d}a\,{\rm d}t$
(17c) $=\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}P_{\tau}(a)\cdot\dot{X}_{\tau}(a)\,p_{0}(a){\rm d}a\,{\rm d}\tau-\int_{\mathbb{R}^{d_{x}}}\left(v_{T}(X_{T}(a))-v_{0}(X_{0}(a))\right)\,p_{0}(a){\rm d}a,$
where we have introduced the momentum variable
(18) $P_{\tau}(a)=\nabla_{x}v_{\tau}(X_{\tau}(a)),$
which amounts to an irrotational fluid in the sense of [27]. Here we have used
(19) $\partial_{t}v_{t}(x=X_{\tau=t}(a))=\partial_{\tau}v_{\tau}(X_{\tau}(a))-\nabla_{x}v_{\tau}(X_{\tau}(a))\cdot\dot{X}_{\tau}(a)$
as well as the fact that the total time derivative $\partial_{\tau}v_{\tau}(X_{\tau}(a))$ contributes boundary terms only. Furthermore,
(20a) $\int_{\mathbb{R}^{d_{x}}}\beta_{t}(x)\cdot\nabla_{x}(\phi_{t}(x)-v_{t}(x))\,\rho_{t}(x){\rm d}x=$
(20b) $\qquad\qquad\int_{\mathbb{R}^{d_{x}}}\tilde{\beta}_{\tau}(a)\cdot\nabla_{x}\left(\phi_{\tau}(X_{\tau}(a))-X_{\tau}(a)\cdot P_{\tau}(a)\right)\,p_{0}(a){\rm d}a$
with $\tilde{\beta}_{\tau}(a)=\beta_{t}(X_{\tau}(a))$.
The equations of motion in $(X_{\tau},P_{\tau})$ arise now from the phase space action principle [4, 9] applied to
(21) $\mathcal{S}(X,P,\beta)=\int_{\mathbb{R}^{d_{x}}}\left\{\int_{0}^{T}\left(P_{\tau}\cdot\dot{X}_{\tau}-H(X_{\tau},P_{\tau},\beta_{\tau})\right){\rm d}\tau-f(X_{T})\right\}p_{0}{\rm d}a$
with Hamiltonian density
(22a) $H(x,p,\beta)(a)$ $:=-\frac{1}{2}\|RG(x(a))^{\rm T}p(a)\|^{2}_{R}+p(a)^{\rm T}b(x(a))+c(x(a))\,+$
(22b) $\qquad\qquad\beta(a)\cdot\nabla_{x}\left(\phi(x(a))-x(a)^{\rm T}p(a)\right)+\frac{1}{2}\Sigma:D_{x}^{2}\phi(x(a)).$
The desired optimal closed loop control law (5) is provided by
(23) $u_{\tau}(X_{\tau}(a))=-RG(X_{\tau}(a))^{\rm T}P_{\tau}(a).$
Upon introducing the generator
(24) $\mathcal{L}_{\beta}g=\beta\cdot\nabla_{x}g+\frac{1}{2}\Sigma:D_{x}^{2}g$
and
(25) $g(x)=\phi(x)-x^{\rm T}p,$
we can rewrite (22) in more compact form as
(26a) $H(x,p,\beta)(a)$ $:=-\frac{1}{2}\|RG(x(a))^{\rm T}p(a)\|^{2}_{R}+p(a)^{\rm T}b(x(a))+c(x(a))\,+$
(26b) $\qquad\qquad\mathcal{L}_{\beta}\left(\phi(x(a))-x(a)^{\rm T}p(a)\right).$
Variation with respect to $\beta_{\tau}$ leads to the constraint
(27) $\nabla_{x}\phi_{\tau}(X_{\tau}(a))=P_{\tau}(a),$
which defines the function $\phi_{\tau}(x)$ in terms of $X_{\tau}(a)$ and $P_{\tau}(a)$ up to a constant. Using the thus specified gradient of $\phi_{\tau}(x)$, we obtain the closed loop control
(28) $u_{\tau}(x)=-RG(x)^{\rm T}\nabla_{x}\phi_{\tau}(x).$
Finally, variations with respect to $X_{\tau}$ and $P_{\tau}$ lead to the Hamiltonian evolution equations
(29a) $\dot{X}_{\tau}(a)$ $=\nabla_{p}H(X_{\tau}(a),P_{\tau}(a),\beta_{\tau}(a)),$
(29b) $-\dot{P}_{\tau}(a)$ $=\nabla_{x}H(X_{\tau}(a),P_{\tau}(a),\beta_{\tau}(a))$
for each $a\in\mathbb{R}^{d_{x}}$. The boundary conditions are $X_{0}(a)=a\sim p_{0}$ and $P_{T}(a)=\nabla_{x}f(X_{T}(a))$. The equations (29) are of mean-field type through $\phi_{t}(x)$, which is determined by the constraint (27).
Dropping labels $a\in\mathbb{R}^{d_{x}}$ from now on and denoting time by $t$ again, the Hamiltonian equations of motion (29) become
(30a) $\dot{X}_{t}$ $=b(X_{t})-G(X_{t})RG(X_{t})^{\rm T}P_{t}-\beta_{t}$
(30b) $=b(X_{t})-G(X_{t})RG(X_{t})^{\rm T}P_{t}-\mathcal{L}_{\beta_{t}}X_{t},$
(30c) $-\dot{P}_{t}$ $=\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}\,+$
(30d) $\qquad\qquad\nabla_{x}\left\{c(X_{t}^{(i)})-\frac{1}{2}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\mathcal{L}_{\beta_{t}}(\phi_{t}(X_{t})-X_{t}\cdot P_{t})\right\}$
(30e) $=\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\nabla_{x}c(X_{t}^{(i)})+\mathcal{L}_{\beta_{t}}\nabla_{x}\phi_{t}(X_{t}),$
which are to be compared to the classical PMP formulation (7). The only new terms arise in connection with the generator $\mathcal{L}_{\beta_{t}}$.
The following theorem provides the key result with regard to the gauge variable $\beta_{t}$ and demonstrates that (30) indeed delivers the desired extension of the PMP to stochastic optimal control problems.
Theorem 3.1.
For any choice of the gauge variable $\beta_{t}$, the resulting function $\phi_{t}(x)$ satisfies
(31) $\nabla_{x}\phi_{t}(x)=\nabla_{x}v_{t}(x)$
where $v_{t}(x)$ is the value function satisfying the HJB equation (6).
Proof.
Let us derive the evolution equation for $\nabla_{x}\phi_{t}(x)$ implied by (27):
(32a) $-\partial_{t}\nabla_{x}\phi_{t}(X_{t})$ $=D_{x}^{2}\phi_{t}(X_{t})\dot{X}_{t}-\dot{P}_{t}$
(32b) $=D_{x}^{2}\phi_{t}(X_{t})\left(b(X_{t})-G(X_{t})R^{-1}G(X_{t})^{\rm T}P_{t}-\beta_{t}\right)\,+$
(32c) $\qquad\frac{1}{2}\nabla_{x}(\Sigma:D_{x}^{2}\phi_{t}(X_{t}))+\nabla_{x}c(X_{t})-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|_{R}^{2}\,+$
(32d) $\qquad\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}+D_{x}^{2}\phi_{t}(X_{t})\beta_{t}$
(32e) $=D_{x}^{2}\phi_{t}(X_{t})\,b(X_{t})-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}\nabla_{x}\phi_{t}(X_{t})\|^{2}_{R}+\nabla_{x}c(X_{t})\,+$
(32f) $\qquad\frac{1}{2}\nabla_{x}(\Sigma:D_{x}^{2}\phi_{t}(X_{t}))+\left(D_{x}b(X_{t})\right)^{\rm T}\nabla_{x}\phi_{t}(X_{t}).$
Hence, since it follows that $\tilde{y}_{t}:=\nabla_{x}\phi_{t}$ satisfies the reverse time partial differential equation
(33) $-\partial_{t}\tilde{y}_{t}=\nabla_{x}(\tilde{y}_{t}\cdot b)+\frac{1}{2}\nabla_{x}(\Sigma:D_{x}\tilde{y}_{t})+\nabla_{x}c-\frac{1}{2}\nabla_{x}\|RG^{\rm T}\tilde{y}_{t}\|^{2}_{R}$
and $\tilde{y}_{T}=\nabla_{x}f$, we find from (8) that $\tilde{y}_{t}=y_{t}$ regardless of the choice of the gauge function $\beta_{t}$. ∎
In line with (15), a natural choice for the gauge function is
(34) $\beta_{t}=\frac{1}{2}\Sigma\nabla_{x}\log\rho_{t}(X_{t}),$
where $\rho_{t}(x)$ denotes the law of $X_{t}$ as defined by (13). This choice leads to
(35) $\mathcal{L}_{\beta_{t}}g=-\frac{1}{2}\Sigma\nabla_{x}\log\rho_{t}\cdot\nabla_{x}g+\frac{1}{2}\Sigma:D_{x}^{2}g,$
which is the generator of a reversible diffusion process with invariant measure $\rho_{t}$ [19]. Since $\mathcal{L}_{\beta_{t}}^{\dagger}\rho_{t}=0$ for the dual operator $\mathcal{L}_{\beta_{t}}^{\dagger}$ [19], this is also the choice of $\beta_{t}$ which makes the variation with respect to $\phi_{t}$ vanish. We recall that this choice of $\beta_{t}$ leads to $\rho_{t}=\pi_{t}$ with $\pi_{t}$ the law of $X_{t}$ arising from the SDE (1).
Alternatively, consider
(36) $\beta_{t}=-G(X_{t})RG(X_{t})^{\rm T}P_{t},$
which eliminates the control from the evolution equation in $X_{t}$ since
(37) $G(X_{t})U_{t}=-G(X_{t})RG(X_{t})^{\rm T}P_{t}=\beta_{t}.$
Hence a single forward and reverse time simulation of
(38a) $\dot{X}_{t}$ $=b(X_{t}),$
(38b) $-\dot{P}_{t}$ $=\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}\nabla_{x}\phi_{t}(X_{t})\|^{2}_{R}\,+$
(38c) $\qquad\qquad\nabla_{x}c(X_{t})+\frac{1}{2}\nabla_{x}(\Sigma:D_{x}^{2}\phi_{t}(X_{t}))$
subject to $X_{0}=a\sim p_{0}$ and $P_{T}(a)=\nabla_{x}f(X_{T}(a))$ solves the stochastic optimal control problem. We note that such a decoupling is possible even when $\Sigma$ is singular or even zero. This is in difference to the classical as well as the stochastic PMP [2].
Both choices for the gauge variable can also be combined into
(39) $\beta_{t}=\frac{1}{2}\Sigma\nabla_{x}\log\rho_{t}(X_{t})-G(X_{t})RG(X_{t})^{\rm T}P_{t}-G(X_{t})u_{t}^{\rm ref}(X_{t}),$
where $u_{t}^{\rm ref}(x)$ is a known reference control; if available. In the context of NMPC [23, 7], such a control could arise from the last receding horizon control window.
4. Discounted cost optimal control problems
We now consider infinite horizon optimal control problems with discounted cost function (3). Such problems can be dealt with by simply introducing the damping term $-\gamma P_{t}$ into the evolution equation for the momentum variable $P_{t}$ [2]. The boundary condition for $P_{T}$ is replaced by the transversality condition
(40) $\lim_{T\to\infty}e^{-\gamma T}P_{T}(a)\cdot X_{T}(a)=0,$
which follows from (21) adapted to the case of discounted cost; i.e.
(41) $\mathcal{S}(X,P,\beta)=\int_{\mathbb{R}^{d_{x}}}\left\{\int_{0}^{\infty}e^{-\gamma t}\left(P_{t}^{\rm T}\dot{X}_{t}-H(X_{t},P_{t},\beta_{t})\right){\rm d}t\right\}p_{0}{\rm d}a,$
and the fact that boundary terms arising from integration by parts should vanish.
A formal derivation of the action function (41) can again start from a reformulation of the associated HJB equation:
(42a) $-\partial_{t}v_{t}$ $=-\gamma v_{t}+b\cdot\nabla_{x}v_{t}+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\phi_{t})+c-\frac{1}{2}\|RG(x)^{\rm T}\nabla_{x}v_{t}\|^{2}_{R},$
(42b) 00 0 $=\phi_{t}-v_{t}.$
The second step is to introduce Lagrange multipliers $\rho_{t}$ and $\beta_{t}$ in order to enforce (42) in terms of the action functional
(43a) $\mathcal{S}(\tilde{v},\rho,\beta)$ $=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\partial_{t}\tilde{v}_{t}+\nabla_{x}\tilde{v}_{t}\cdot b+e^{-\gamma t}c-\frac{e^{\gamma t}}{2}\|RG^{\rm T}\nabla_{x}\tilde{v}_{t}\|^{2}_{R}\right\}\rho_{t}{\rm d}x\,{\rm d}t\,+$
(43b) $\qquad\qquad\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\beta_{t}\cdot\nabla_{x}(\tilde{\phi}_{t}-\tilde{v}_{t})+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\tilde{\phi}_{t})\right\}\rho_{t}{\rm d}x\,{\rm d}t.$
Here we have scaled the value function by $e^{-\gamma t}$; i.e. $\tilde{v}_{t}=e^{-\gamma t}v_{t}$. After transferring to the Lagrangian framework, we undo the scaling in $\tilde{P}_{t}=\nabla_{x}\tilde{v}_{t}$ to obtain the action functional (41).
Since we are interested in equilibrium states only; i.e. $t\to\infty$, we can run both equations forward in time resulting in
(44a) $\dot{X}_{t}$ $=\nabla_{p}H(X_{t},P_{t},\beta_{t}),$
(44b) $\dot{P}_{t}$ $=-\gamma P_{t}+\nabla_{x}H(X_{t},P_{t},\beta_{t}).$
Then $(X_{\infty}(a),P_{\infty}(a))$ define the optimal control law (28) via the implied $\nabla_{x}\phi_{\infty}(x)$. In this setting, the choice (34) for $\beta_{t}$ emerges as the most natural one.
We note that such a joint forward integration is not possible for forward-backward SDE formulations of the PMP [2].
5. Numerical implementation
Numerical implementations of (30) rely on a set of particles $(X_{t}^{(i)},P_{t}^{(i)})$, $i=1,\ldots,M$, which evolve according to
(45a) $\dot{X}_{t}^{(i)}$ $=\nabla_{P_{i}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\}),$
(45b) $-\dot{P}_{t}^{(i)}$ $=\nabla_{X_{i}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\})$
for appropriate Hamiltonian $\mathcal{H}$ and subject to boundary conditions
(46) $X_{0}^{(i)}=a_{i}\sim\pi_{0},\qquad P_{T}^{(i)}=-\nabla_{x}f(X_{T}^{(i)}).$
In order to get closed form ODEs, we need to specify $\mathcal{H}$ which involves approximating $\phi_{t}(x)$ as defined by the constraint (27). We discuss a few choices below. These approximations apply equally well to the discounted cost mean-field PMP formulation from Section 4.
5.1. A regression formulation
The constraint (27) can be approximated via the method of least squares. For example, one could use
(47) $\mathcal{V}_{t}(\gamma_{t})=\frac{1}{2M}\sum_{i=1}^{M}\|P_{t}^{(i)}+\nabla_{x}\tilde{\phi}(X_{t}^{(i)};\gamma_{t})\|^{2}$
for suitable class of functions $\tilde{\phi}(\cdot;\gamma)$ parametrized by $\gamma\in\mathbb{R}^{d_{\gamma}}$. This approach is not strictly variational unless $\Sigma=0$; but it can be applied to any choice of the gauge variable $\beta_{t}(a)$. See Section 5.3 below for a strictly variational approach.
As a specific example, consider
(48) $\beta_{t}^{(i)}=\frac{1}{2}G(X_{t}^{(i)})R^{-1}G(X_{t}^{(i)})(A_{t}X_{t}^{(i)}+c_{t}-P_{t}^{(i)})$
together with a bilinear ansatz
(49) $\tilde{\phi}(x;\gamma)=\frac{1}{2}x^{\rm T}Ax+x^{\rm T}c,$
which makes the term depending on $\Sigma$ in (22) vanish. Thus the Hamiltonian (22) gives rise to
(50a) $\mathcal{H}_{t}(\{X_{t}^{(j)}\},\{P_{t}^{(i)}\})$ $=\sum_{i=1}^{M}\left\{-\frac{1}{2}\|RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}\|^{2}_{R}+(P_{t}^{(i)})^{\rm T}b(X_{t}^{(i)})+c(X_{t}^{(i)})\,+\right.$
(50b) $\qquad\qquad\left.\frac{1}{2}\|RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}-c_{t}-P_{t}^{(i)})\|^{2}_{R}\right\}$
and the equations of motion (45) reduce to
(51a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}\left(A_{t}X_{t}^{(i)}+c_{t}\right),$
(51b) $-\dot{P}_{t}^{(i)}$ $=(D_{x}b(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})+\frac{1}{2}\nabla_{x}\|RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}+c_{t})\|_{R}^{2}\,-$
(51c) $\qquad\qquad\nabla_{x}(G(X_{t}^{(i)})P_{t}^{(i)})^{\rm T}RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}+c_{t})$
with $A_{t}$ and $c_{t}$ determined by
(52) $A_{t}=C_{t}^{px}(C_{t}^{xx})^{-1},\qquad c_{t}=m^{p}_{t}-A_{t}m^{x}_{t}.$
Here $C_{t}^{px}$ and $C_{t}^{xx}$ denote the empirical covariance matrices between $\{P_{t}^{(i)}\}$ and $\{X_{t}^{(i)}\}$, and $\{X_{t}^{(i)}\}$ with itself, respectively. The empirical mean of $\{X_{t}^{(i)}\}$ and $\{P_{t}^{(i)}\}$, respectively, is denoted by $m_{t}^{x}$ and $m_{t}^{p}$, respectively. Note that $A_{t}$ is not symmetric in general; but can be made symmetric via $1/2(A_{t}^{\rm T}+A_{t})$.
Formulation (51) simplifies further in case $G$ does not depend on $x$; i.e.
(53a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-GRG^{\rm T}\left(A_{t}X_{t}^{(i)}+c_{t}\right),$
(53b) $-\dot{P}_{t}^{(i)}$ $=(D_{x}b(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})+A^{\rm T}G^{\rm T}RG^{\rm T}(A_{t}X_{t}^{(i)}+c_{t}-P_{t}^{(i)}).$
While the regression approach can be combined with any choice of $\beta_{t}$, the following two approaches rely on the specific choice (34).
5.2. Discrete Schrödinger bridge approximation
Let us consider (26) with $\beta_{t}$ given by (34) in more detail. We replace the resulting generator (35) by a discrete Schrödinger bridge approximation $m_{t}^{(ij)}$ [16, 28, 8] over the set of particles $\{X_{t}^{(j)}\}$ giving rise to
(54a) ($\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\})$ $=\frac{1}{M}\sum_{i=1}^{M}\left\{-\frac{1}{2}\|RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}\|_{R}^{2}+P_{t}^{(i)}\cdot b(X_{t}^{(i)})\right\}\,+$
(54b) $\qquad\qquad\sum_{i=1}^{M}\left\{G_{t}^{(i)}(\{X_{t}^{(j)}\})+\beta_{t}^{(i)}\cdot P_{t}^{(i)}+c(X_{t}^{(i)})\right\}$
with
(55) $\beta_{t}^{(i)}:=-\sum_{i=1}^{M}m_{t}^{(ij)}X_{t}^{(j)}$
and
(56) $G_{t}^{(i)}(\{X_{t}^{(j)}\}):=\sum_{j=1}^{M}m_{t}^{(ij)}\phi_{t}(X_{t}^{(j)}).$
The coefficients $m_{t}^{(ij)}\geq 0$ are defined by
(57) $m_{t}^{(ij)}=\frac{1}{\varepsilon}\left(v_{t}^{(i)}d_{t}^{(ij)}v_{t}^{(j)}-\delta_{ij}\right),\qquad d_{t}^{(ij)}=e^{-\frac{1}{2\varepsilon}\|X_{t}^{(i)}-X_{t}^{(j)}\|_{\Sigma}^{2}},$
with $\{v_{t}^{(i)}\}$ chosen such that
(58) $\sum_{i}m_{t}^{(ij)}=\sum_{j}m_{t}^{(ij)}=0.$
Here $\delta_{ij}$ denotes the Kronecker delta function; i.e., $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ otherwise. The regularization parameter $\varepsilon>0$ needs to be chosen appropriately [28]. A particularly elegant formulation arises from the choice $\varepsilon=\Delta t$ with $\Delta t>0$ the step-size of an Euler discretization of the Hamiltonian equations of motion [8]. A fast iteterative algorithm for computing the coefficients $v_{t}^{(i)}\geq 0$ has been given in [16, 28].
Writing down the associated Hamiltonian equations of motion (45), we obtain a fully variational implementation of the proposed PMP. However, computing gradients of $m_{t}^{(ij)}$ with respect to particle positions $X_{t}^{(k)}$ is non-trivial and we use the simpler approximation
(59a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}-\sum_{j=1}^{M}m_{t}^{(ij}X_{t}^{(j)},$
(59b) $-\dot{P}_{t}^{(i)}$ $=(Db(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\sum_{j=1}^{M}m_{t}^{(ij)}P_{t}^{(j)}.$
The discrete Schrödinger bridge approximation does not require an explicit approximation of $\phi_{t}(x)$, which simplifies numerical implementations. If $\nabla_{x}\phi_{t}(x)$ is needed it can be obtained via (47).
5.3. Variational approach
A particularly versatile approach is to approximate $\beta_{t}$ and $\phi_{t}$ via a parametric ansatz $\beta_{t}(x)=\tilde{\beta}(x;\gamma)$ and $\phi_{t}(x)=\tilde{\phi}(x;\alpha_{t})$, respectively, and to consider the empirical Hamiltonian
(60a) ($\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma,\alpha)\,=$
(60b) $\qquad\qquad\sum_{j=1}^{M}\left\{-\frac{1}{2}\|RG(X_{t}^{(j)})^{\rm T}P_{t}^{(j)}\|^{2}_{R}+P_{t}^{(j)}\cdot b(X_{t}^{(j)})+c(X_{t}^{(j)})\,+\right.$
(60c) $\qquad\qquad\left.\tilde{\beta}_{t}(X_{t}^{(j)},\gamma)\cdot\nabla_{x}\left(\tilde{\phi}(X_{t}^{(j)},\alpha)-X_{t}^{(j)}\cdot P_{t}^{(j)}\right)+\frac{1}{2}\Sigma:D_{x}^{2}\tilde{\phi}(X_{t}^{(j)};\alpha)\right\}$
with equations of motion
(61a) $\dot{X}_{t}^{(i)}$ $=\nabla_{X^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma_{t},\alpha_{t}),$
(61b) $-\dot{P}_{t}^{(i)}$ =($=\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma_{t},\alpha_{t}),$
(61c) $(\gamma_{t},\alpha_{t})$ $=\arg\min_{(\gamma,\alpha)}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma,\alpha).$
Appropriate choices for $\tilde{\beta}$ and $\tilde{\phi}$ can follow the approaches considered, for exmaple, in [15].
6. Numerical experiments
We present results from two numerical examples. Both cases involve infinite horizon cost functions of the form (3) and controlled ODEs. We employ the initial value formulation (44) and its numerical approximation using discrete Schrödinger bridges as outlined in Section 5.2. In both cases, we add a diffusion (viscosity) regularization in (22) in line with the concept of viscosity solutions to the associated HJB equations and the vanishing viscosity method [3].
6.1. Inverted pendulum
Figure 1. Upper right panel: Time evolution of the ensemble mean angle and ensemble mean velocity. Upper left panel: Control law (65) at final time as function of $(\theta,v)$. Lower right panel: Control law (65) at final time as function of $\theta$. Lower left panel: Control law (65) at final time as function of $v$.
We consider a controlled inverted pendulum with friction [18]. The state variable $X=(\theta,v)^{\rm T}$ is two-dimensional with equations of motion
(62a) $\dot{\theta}_{t}$ $=v_{t},$
(62b) $\dot{v}_{t}$ $=\sin(\theta_{t})-\sigma v_{t}+\cos(\theta_{t})U_{t}$
and $\sigma=5$. Consider the running cost function
(63) $c(x)=\frac{\alpha}{2}\left(\theta^{2}+v^{2}\right)$
with $\alpha=25$ and set $R=1$ in (4). We use the discount factor $\gamma=1$ in (3). Note that $(0,0)$ is an unstable equilibrium point under $U_{t}\equiv 0$.
In order to implement and, in fact, regularize the mean-field PMP, we add diffusion (viscosity) by setting $\Sigma=0.1I$ in (22). We initialize the particles at the stable equilibrium point $x_{\rm s}=(\pi,0)$ with added Gaussian noise of variance $0.1I$; i.e. $X_{0}^{(i)}\sim{\rm N}(x_{\rm s},0.1I)$. The initial momenta are set to $P_{0}^{(i)}=-2SX_{0}^{(i)}$ with
(64) $S=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right).$
The discrete Schrödinger bridge based formulation from Section 5.2 has been implemented with ensemble size $M=200$ and $\varepsilon=\Delta t$. The resulting evolution equations (44) have been simulated over 8000 Euler time-steps with time-step $\Delta t=0.025$.
The results can be found in Figure 1. It can be clearly seen that the mean-field formulation leaves the stable equilibrium and rapidly equilibrates at the unstable equilibrium. The resulting final control terms
[…]
strongly depend on $\theta$ while the dependence on $v$ is less structured.
We also note that keeping a non-zero diffusion matrix $\Sigma=0.1I$ in (22) leads to a desirable exploration of state space and, hence, to a stabilization/regularization of the control problem. However, a wider exploration of state space implies the need for larger ensemble sizes. Hence a balance of computational cost and state space exploration via added diffusion is required in practice and needs to be explored further.
6.2. Lorenz-63
Figure 2. Controlled Lorenz-63 model: Displayed is the ${\rm x}$-component of a single particle trajectory chosen at random from the ensemble $\{X_{t}^{(i)}\}$ and its associated control $U_{t}^{(i)}$ as a function of time.
We follow [12] and consider an optimal control problem for the Lorenz-63 model
(66) $\dot{X}_{t}=f(X_{t})+U_{t},\qquad f(X)=\left(\begin{array}[]{l}\sigma({\rm y}-{\rm x})\\ -{\rm x}{\rm z}+r{\rm x}-{\rm y}\\ {\rm x}{\rm y}-b{\rm z}\end{array}\right),$
in the state variable $X=({\rm x},{\rm y},{\rm z})^{\rm T}\in\mathbb{R}^{3}$ and with parameters $\sigma=10$, $r=28$, and $b=8/3$ [13]. The control problem is to restrict solutions of (66) to ${\rm x}_{t}\geq 0$ and the authors of [12] introduce the running cost
(67) $c(X)=\frac{\alpha}{2}\left(\min({\rm x},0)\right)^{2}.$
In our experiments, we have followed this cost function and have set $\alpha=5000$ and $R=1$ in (4). Contrary to the setting of [12], we have implemented a discounted cost function (3) with $\gamma=1$ instead of a receding horizon NMPC approach [7, 23].
We have implemented the Schrödinger bridge based mean-field PMP formulation with $\varepsilon=0.1$, step-size $\Delta t=0.001$, and $M=100$ particles. We again added artificial diffusion (viscosity) via $\Sigma=0.5I$ in (22). The results can be found in Figures 2 and 3. The impact of the control can be clearly detected. In particular, by enforcing ${\rm x}_{t}\geq 0$, the chaotic nature of the Lorenz-63 model has been eliminated. The initial spiral towards to quasi-periodic attractor is due to the chosen initialization
(68) $X_{0}=(7.8590;7.1136;27.2293)^{\rm T}+\Xi_{0},\qquad\Xi_{0}\sim{\rm N}(0,0.1I),$
of the Lorenz-63 problem. Our results are in qualitative agreement with the findings from [12] which are based on a combination of data assimilation [26, 6] and NMPC [7, 23].
Figure 3. Controlled Lorenz-63 model: Displayed is the three-dimensional trajectory of a single particle chosen at random from the controlled ensemble $\{X_{t}^{(i)}\}$.
7. Conclusions
A deterministic mean-field formulation of the PMP has been presented. Similar to the well-known forward-backward SDE formulation of the stochastic PMP [2] and its numerical implementation; see e.g. [5], a regression problem (27) arises when discretized as an interacting particle system. Several numerical implementations have been discussed. Contrary to its stochastic counterpart, the mean-field approach leads to smooth particle trajectories which should help with reducing the variance in the involved approximations. Furthermore, the freedom in the choice of the gauge variable $\beta_{t}$ allows for a decoupling of the forward and backward ODEs in $X_{t}$ and $P_{t}$, respectively. Furthermore, a forward simulation of the coupled Hamiltonian ODEs is possible in case of infinite horizon optimal control problems. Numerical results have been presented for a controlled inverted pendulum as well as a controlled Lorenz-63 system. While the numerical experiments focus on infinite horizon control problems and the Schrödinger bridge approximation, application of the proposed methodology to NMPC [23, 7] and variational approximations will be investigated in future work. Another area of interest is provided by extensions to partially observed optimal control problems [22, 1, 17, 29, 24].
Acknowledgment. This work has been funded by Deutsche Forschungsgemeinschaft (DFG) - Project-ID 318763901 - SFB1294.